OpenStudy (anonymous):

Hoh to find exact value of sin(3) (degress) ?

OpenStudy (anonymous):

Can you use a calculator or not??

OpenStudy (anonymous):

not use calcul, but use trignometry concept..

OpenStudy (unklerhaukus):

can you find the sine of 12°, 15°

OpenStudy (anonymous):

no, i can't..how to get it, please give a example

OpenStudy (unklerhaukus):

\[\sin(u\pm v)=\sin (u) \cos(v)\pm \cos(u)\sin(v)\]

OpenStudy (unklerhaukus):

\[\sin(3°)=\sin(15°- 12°)=\sin (15°) \cos(12°)- \cos(15°)\sin(12°)\]

OpenStudy (anonymous):

yea, i see..so, how to calculate sin15 and sin12 ??

OpenStudy (anonymous):

Uncle Rocks I think we have to go with 18 and 15.. It is very difficult to calculate sin12 and cos12 I think.. Or you have a method to find sin12???

OpenStudy (unklerhaukus):

i havent worked this through, i was just thinking that 12/360=1/30, 18 might be easier, as 18/360=1/20

OpenStudy (anonymous):

I have the value of sin and cos 15 and 18, that is why I am thinking to use that..

OpenStudy (unklerhaukus):

go for it waterineyes !

OpenStudy (anonymous):

Should I give the values directly or I have to derive them fully here how they come ??

OpenStudy (unklerhaukus):

derive them from stuff we should know

OpenStudy (anonymous):

Firstly, Let us find sin18.. let x = 18 5x = 90 2x + 3x = 90 2x = 90 - 3x take sin both sides, sin2x = sin(90-3x) as sin(90-y) = cosy So, sin2x = cos3x Here we have to use Identities; \[\sin2x = 2sinx.cosx\] \[\cos3x = 4\cos^3x - 3cosx\] So, \[2sinx.cosx = 4\cos^3x - 3cosx\] Bringing on one side and taking cosx common, \[cosx(2sinx - 4\cos^2x + 3) = 0\] \[2sinx - 4(1-\sin^2x) + 3 = 0\] \[4\sin^2x + 2sinx - 1 = 0\] Solving this by quadratic formula we get: \[sinx = \frac{-2 \pm \sqrt{20}}{8}\] or, \[sinx = \frac{-1 \pm \sqrt{5}}{4}\] One point to be noted here that the x = 18 lies in 1st quadrant so value of sinx will be positive here,. So, \[sinx = \frac{-1 +\sqrt{5}}{4} = \frac{\sqrt{5} - 1}{4}\]

OpenStudy (anonymous):

Now we know that: \[cosx = \sqrt{1 - \sin^2x}\] By this we get the value of: \[\cos18 = \frac{\sqrt{10 + 2\sqrt{5}}}{4}\]

OpenStudy (anonymous):

It is getting more complex now..

OpenStudy (anonymous):

values of sin15 and cos15 are easier to find though.. \[\sin(45-30) = \sin45\cos30 - \sin30 \cos45\] \[\cos(45 - 30) = \cos45. \cos30 + \sin45.\sin30\]

OpenStudy (anonymous):

After solving we get, \[\sin15 = \frac{\sqrt{3} - 1}{2\sqrt{2}}\] \[\cos15 = \frac{\sqrt{3} + 1}{2\sqrt{2}}\]

OpenStudy (anonymous):

Now, lots of work is remaining to find sin3, \[\sin(18-15) = \sin18.\cos15 - \sin15.\cos18\]

OpenStudy (anonymous):

waawwww, very very very long time to finished... amzing,,, i think i will giveus if i do the problem, but i like your solution,, thank you for all... BTW, from sin3, we can find the value sin 1 and sin 2 isn't it?? hehe

OpenStudy (anonymous):

Ha ha ha.. I will definitely use a calculator for this.. Ha ha..

OpenStudy (unklerhaukus):

do can you find \(\sin1°, \sin2°\) @tanjung ?

OpenStudy (unklerhaukus):

how can you*

OpenStudy (anonymous):

hahahaha,,, very simple i hope i can do it

OpenStudy (anonymous):

I think he or she is joking.. Isn't it?? @tanjung

OpenStudy (anonymous):

sinx= x- x^3/3!+ x^5/5! - x^7/ 7!+ ...... so on put 3 and get the value

OpenStudy (anonymous):

what is it, Taylor series??

OpenStudy (anonymous):

using complex numbers ,,you can get this expression... cosx= x- x^2/2! +x^4/4!.... so on